OpenStudy (phebe):

12. Using the discriminant, how many real number solutions does this equation have? 3x2 – 2 = 5x @ShadowLegendX

OpenStudy (nincompoop):

do you know the discriminant?

OpenStudy (phebe):

no

OpenStudy (nincompoop):

you need to arrange your equation into proper format first Ax^2 + Bx + C = 0

OpenStudy (phebe):

ok

OpenStudy (phebe):

i did that

OpenStudy (nincompoop):

write it down here so I can see if you did it correctly

OpenStudy (phebe):

ok hold on

OpenStudy (phebe):

um.. jus a sec

OpenStudy (phebe):

do i substutute

OpenStudy (nincompoop):

substitute what?

OpenStudy (phebe):

ohhh nvm im not suppose to do that

OpenStudy (phebe):

do i use PEMDAS

OpenStudy (nincompoop):

turn your equation into zero \[3x^2-5x-2 = 0\]

OpenStudy (phebe):

okk want me to do it lik that

OpenStudy (phebe):

so is 3x^2-5x-2=0 the answer

OpenStudy (nincompoop):

not yet it is just the beginning of figuring out the discriminant

OpenStudy (phebe):

ok

OpenStudy (nincompoop):

do you know the quadratic formula?

OpenStudy (phebe):

yess kinda

OpenStudy (nincompoop):

start identifying which is A, B and C these are the coefficients of the equation

OpenStudy (phebe):

ok

OpenStudy (nincompoop):

do you know what coefficients are?

OpenStudy (phebe):

yess

OpenStudy (phebe):

do u know the answer

OpenStudy (phebe):

is the answer 2

OpenStudy (phebe):

@mathstudent55

OpenStudy (nincompoop):

show me the solution how you got 2

OpenStudy (phebe):

i jus used the link u showed me thas all.

OpenStudy (phebe):

I got 2 because its a coefficient

OpenStudy (phebe):

wait coeffients are lik the ones with no exopnent or no

OpenStudy (nincompoop):

A = 3 B = -5 C = -2 coefficients are the number in front of the variables in an expression.

OpenStudy (phebe):

ok i got that

OpenStudy (phebe):

so wats the answer

OpenStudy (nincompoop):

use the discriminant formula I provided

OpenStudy (phebe):

ok well i got 2 is that right

OpenStudy (nincompoop):

can you write it down and show me how you got 2

OpenStudy (phebe):

well i jus guessed

OpenStudy (phebe):

am i right

OpenStudy (nincompoop):

no

OpenStudy (phebe):

ohhh lol well can u tel me the answer

OpenStudy (nincompoop):

b^2 -4ac (-5)^2 - 4(3 * -2)

OpenStudy (phebe):

thas the answer

OpenStudy (phebe):

thanx @nincompoop nd fan me lol u neva did

OpenStudy (anonymous):

poor phoebe